OpenStudy (anonymous):

12 * (-5/12) Show work if you can please :D

OpenStudy (anonymous):

\[12\div \left( -5/12 \right) =???\]

OpenStudy (anonymous):

the 12 is like a 12/1 so it cancels the 12 in the denominator o fthe -5\[12*\frac{-5}{12}=-5\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Thank you Hehe:D

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

7 1/5 *(-1 3/10)+???